Tire pressure monitoring systems (TPMS) warn the driver when the tire pressure of the vehicle is 25% below the target pressure. Suppose the target tire pressure of a certain car is 32 psi (pounds per square inch.)

(a) At what psi will the TPMS trigger a warning for this car? (Round your answer to 2 decimal place.)

When the tire pressure is __BELOW__24 psi.

(b) Suppose tire pressure is a normally distributed random variable with a standard deviation equal to 3 psi. If the car’s average tire pressure is on target, what is the probability that the TPMS will trigger a warning? (Round your answer to 4 decimal places.)

Probability

(c) The manufacturer’s recommended correct inflation range is 30 psi to 34 psi. Assume the tires’ average psi is on target. If a tire on the car is inspected at random, what is the probability that the tire’s inflation is within the recommended range? (Round your intermediate calculations and final answer to 4 decimal places.)

Probability
Answer #1

a)

When the tire pressure is BELOW =32*(1-0.25)=24

b)

probability =P(X<24)=(Z<(24-32)/3)=P(Z<-2.67)=0.0038

c)

probability =P(30<X<34)=P((30-32)/3)<Z<(34-32)/3)=P(-0.67<Z<0.67)=0.7486-0.2514=0.4972
Answer #2

(a) The TPMS triggers a warning when the tire pressure is 25% below the target pressure of 32 psi.

25% of 32 psi = 0.25 * 32 psi = 8 psi

The TPMS will trigger a warning when the tire pressure is 32 psi - 8 psi = 24 psi.

(b) Let X be the tire pressure, which is normally distributed with a mean (average) of 32 psi and a standard deviation of 3 psi.

To find the probability that the TPMS will trigger a warning, we need to find the probability that X is less than 24 psi.

Z-score = (X - mean) / standard deviation Z = (24 - 32) / 3 = -2.67 (rounded to 2 decimal places)

Now, we need to find the probability that Z is less than -2.67 using a standard normal distribution table or a calculator.

Probability (Z < -2.67) ≈ 0.0038 (rounded to 4 decimal places)

Therefore, the probability that the TPMS will trigger a warning when the car's average tire pressure is on target is approximately 0.0038 or 0.38%.

(c) To find the probability that a tire's inflation is within the recommended range of 30 psi to 34 psi, we need to find the probability that X (tire pressure) falls between 30 psi and 34 psi.

Z-score for 30 psi: Z1 = (30 - 32) / 3 = -0.67 (rounded to 2 decimal places)

Z-score for 34 psi: Z2 = (34 - 32) / 3 = 0.67 (rounded to 2 decimal places)

Now, we need to find the probability that -0.67 < Z < 0.67 using a standard normal distribution table or a calculator.

Probability (-0.67 < Z < 0.67) ≈ 0.5040 (rounded to 4 decimal places)

Therefore, the probability that a tire's inflation is within the recommended range of 30 psi to 34 psi is approximately 0.5040 or 50.40%.


answered by: mervetokaz